When $ 250 added to 1/4th of a given amount of money makes it smaller than 1/3 rd of the given amount of money by $ 100. What is the given amount of money?
(a) $ 350 (b) $ 600 (c) $ 4200 (d) $ 3600

Answer by mananth(16949)

Let x be the amount distributed
x/4 +250 = x/3-100
x/4 - x/3 = -250-100
3x-4x /12 = -350
-x= -350*12
x=4200
the distributed amoount = $4200